# Shedding Light on Task Decomposition in Program Synthesis: The Driving Force of the Synthesizer Model

### v1

@@ 1 Introduction
ExeDec (Shi et al., 2023b) incorporates these two ideas by combining program synthesis with task decomposition. As such, ExeDec can be seen as a multi-level transductive-inductive reasoning approach, which has shown to be promising for boosting program synthesis performance (Li et al., 2024). It employs two interconnected components: a Subgoal Model that iteratively divides complex tasks into subtasks by predicting the subgoal of the next step (high-level transductive reasoning) and a synthesis model that generates programs to solve these subtasks (low-level inductive reasoning). This structured framework enables the Subgoal Model to acquire broadly applicable skills essential for addressing diverse and complex tasks, while the synthesis model focuses on their implementation.
@@ 3.2 REGISM
To further investigate the role of ExeDec’s Subgoal Model, we developed Repeated Execution-Guided Invoking of Synthesis Model (REGISM), a modified version of the ExeDec framework that excludes the Subgoal Model. REGISM invokes the Synthesizer Model to generate a single-step program, i.e., a program consisting of one function only, executes this program and updates the task specifications using the execution output. Consequently, the REGISM architecture and workflow exactly mirror ExeDec’s No-Subgoal Ablation. The only difference lies in the training paradigm and, thus, the role of the synthesis model. The No-Subgoal Ablation model requires the synthesis model to predict subprograms using the (updated) task inputs and the task outputs, meaning the model must implicitly learn both decomposition strategies and program generation abilities. Since it does not explicitly rely on a separate decomposition module, it must infer subgoals while simultaneously solving the task, making its learning process more entangled. In contrast, the Synthesizer Model used in REGISM is trained to predict subprograms using subtask inputs and subtask outputs instead of the final task output. This allows the model to focus solely on program generation without the added complexity of implicitly learning decomposition.
@@ 3.3.1 Domains
String Manipulation The Robustfill domain, introduced by Devlin et al. (2017), focuses on synthesizing sequences of string manipulation operations. In this domain, I/O pairs consist of strings, and the goal is to generate programs that transform input strings into corresponding output strings. Key features of the domain include the use of a DSL with functions that perform various string operations (Appendix B). These operations include extracting substrings from the input, modifying strings, and composing multiple operations by applying modifications to the results of prior expressions. The DSL also supports the inclusion of constant string characters. Robustfill programs are structured as concatenations of expressions that are largely independent of each other, with the exception of the Compose operation which applies an operation to the result of another function. This weak dependence among subtasks results in a smaller combinatorial space. With fewer viable decompositions available, learning task decomposition becomes more straightforward. However, this independence also limits the ability to perform error correction during synthesis, making it essential to accurately decompose and implement each subtask to ensure correct program generation.
@@ 4.1 Domain Differences and Quality of the Subgoal Model
To test our hypotheses, we conduct a detailed analysis of ExeDec’s performance, focusing on how effectively it learns task decompositions compared to the ground truth. Figure 3 presents a density plot illustrating the distribution of solved tasks.
Figure 3: Tasks solved by ExeDec, clustered based on subtask and subprogram accuracy. The x-axis represents subtask accuracy per task, measured as the overlap between predicted subtasks and ground truth subtasks. The y-axis indicates subprogram accuracy, defined as the overlap between predicted programs and the ground truth solutions. Consequently, tasks solved as intended by the ground truth are clustered in the top-right corner. The values represent averages across all compositional generalization categories. The distributions were computed using over 8,900 solved tasks in the DeepCoder domain and more than 26,000 in the RobustFill domain.
The x-axis represents the subtask accuracy per task, measured as the overlap between predicted subtasks and ground truth subtasks. The y-axis indicates subprogram accuracy, defined as the overlap between predicted programs and the ground truth solutions. For a task that requires four decomposition steps to be solved, if ExeDec predicts two subtask specifications correctly as per the ground truth solution and successfully solves three subtasks with programs matching the ground truth, the corresponding data point for this task would be positioned at ($x=50\%$, $y=75\%$). Consequently, the density plot reveals clusters that capture the semantic and syntactic accuracy of task solutions. In the top-right quadrant, both subtasks and subprograms closely align with the ground truth solution. The bottom-right region represents cases where subtasks mostly match the ground truth, but subprograms differ due to alternative syntactic implementations. The bottom-left quadrant contains instances where both subtasks and subprograms deviate from the ground truth. Finally, in the top-left region, subprograms align closely with the ground truth while subtasks differ.
@@ 4.2 REGISM as a Driving Force in Boosting Performance
REGISM performs close to ExeDec on average To assess the effect of repeated invoking of the Synthesizer Model, we evaluate REGISM on the same tasks as ExeDec. As shown in Figure 5, REGISM performs better than the baseline Synthesizer Model, demonstrating the value of repeated invocation and execution guidance. ExeDec shows improved performance on the training distribution compared to REGISM. Moreover, it also outperforms REGISM in tasks requiring the composition of different concepts, benefiting from its decomposition approach that helps identify and combine relevant subcomponents. REGISM struggles with length generalization tasks, which may result from its training paradigm focusing on single-step tasks. This approach makes deriving correct subprograms from tasks requiring more decompositions more challenging compared to training methods that include multi-step tasks, such as the No-Subgoal ablation. However, excluding length generalization tasks shows that REGISM performs only slightly worse than ExeDec on average, suggesting that repeated invocation of the Synthesizer Model plays an important role. While ExeDec’s decomposition strategies are advantageous in some task categories, they provide limited or no benefit in others and can even hinder performance in certain cases. As REGISM performs similarly to ExeDec and its No-Subgoal Ablation, it appears to be the key factor in solving tasks, even though REGISM does not rely on explicit or implicit decomposition guidance.
@@
Comparing REGISM’s results to ExeDec (Figure 4), the tasks solved by REGISM proof to be of similar difficulty to those solved by ExeDec because the number of ground truth decompositions is comparable. Moreover, REGISM solves tasks in about as many iterations as decompositions are intended by the ground truth, i.e., in significantly fewer steps than ExeDec. This indicates that REGISM solves tasks closer to the ground truth solution thus raising the hypothesis that REGISM’s inherent decompositions are closer to the ground truth decomposition than the explicit ones of ExeDec. To investigate this, we plot the subprogram accuracy against the overlap with the ground truth decomposition after executing the subprograms generated by REGISM (Figure 7). This analysis is further split by tasks solved by both approaches or exclusively by either approach. Our findings reveal several notable patterns: REGISM’s task solutions tend to spread along the main diagonal. There is one prominent cluster at the right edge of each plot, where all subtasks align with the ground truth decomposition. To a similar but lesser extent than in ExeDec, REGISM’s largest cluster is located in the bottom-left quadrant. However, REGISM’s solutions exhibit a more uniform distribution along the x-axis compared to ExeDec, whose distribution is distinctly shifted toward lower subtask accuracies. This effect is slightly more pronounced for tasks solved exclusively by REGISM than for those solved by both approaches. These findings indicate that the Synthesizer Model within REGISM generates solutions closer to the ground truth than ExeDec despite not being explicitly or implicitly trained for this purpose. However, for both approaches, the quality of decompositions does not differ significantly between tasks solved exclusively and those solved by both methods. This finding leaves open the question of why ExeDec solves certain tasks that REGISM does not, especially when the predicted subgoals in those tasks are not more precise than in tasks solved by REGISM, and vice versa.
@@ 5 Related Work
Programming-by-Example Neural networks are extensively used in guiding program synthesis for PBE (Balog et al., 2016; Yin & Neubig, 2017; Lee et al., 2018). Many approaches encompass techniques that divide the synthesis process into multiple steps, facilitating either iterative refinement or task planning. These approaches commonly employ search strategies, including top-down methods that recursively construct programs by expanding partial solutions (Nye et al., 2019), and bottom-up methods that focus on solving smaller problems and combining them into complete programs (Shi et al., 2022; 2023a; Odena et al., 2020). Planning-based methods first generate high-level abstract representations of a program’s structure, which are subsequently refined into executable code (Murali et al., 2017; Nye et al., 2019; Chen et al., 2020; Hong et al., 2021; Klinger et al., 2023; Prasad et al., 2023; Zhang et al., 2023; Cano et al., 2023). Execution-guided approaches, on the other hand, use intermediate feedback, such as partial program executions or state traces, to iteratively adjust and improve programs (Ellis et al., 2019; Chen et al., 2018; Shrivastava et al., 2021). Recent advancements in Large-Language Models have further contributed to program synthesis by utilizing pre-trained models trained on large-scale corpora of code and natural language (Li & Ellis, 2024; Li et al., 2024; Tang et al., 2024). These models are integrated into synthesis frameworks to generate or guide program solutions by leveraging their ability to capture both semantic and structural patterns of code.
Decomposition-based Program Synthesis Multi-step synthesis methods inherently decompose tasks into subtasks as part of their iterative process. In contrast, decomposition approaches aim to (explicitly) decompose a task before exploiting this structure to improve synthesis. Various approaches explicitly model relationships between subprograms, enabling relational reasoning to identify dependencies and shared structures across subtasks (Hocquette & Cropper, 2023; Hocquette et al., 2024; Hocquette & Cropper, 2024). Similarly, hierarchical composition methods explicitly structure programs by progressively building higher-level abstractions from lower-level components, ensuring a more systematic decomposition (Liu et al., 2023).
@@ 6 Conclusion
Hybrid approaches that integrate transductive and inductive solution strategies have shown promise in enhancing program synthesis performance (Li et al., 2024). ExeDec can be seen as such a hybrid model combining both ideas: the Subgoal Model represents the transduction component, while the Synthesizer Model embodies the induction aspect, working together to achieve compositional generalization. Our results demonstrate that both REGISM and ExeDec successfully tackle a variety of tasks. As such, framing ExeDec as a hybrid approach may offer additional advantages. However, the precise circumstances under which this combination proves most effective remain an open area for further exploration.

### v3

@@ 1 Introduction
ExeDec (Shi et al., 2023b) incorporates these two ideas by combining program synthesis with task decomposition. As such, ExeDec can be seen as a multi-level transductive-inductive reasoning approach, which has shown to be promising for boosting program synthesis performance (Li et al., 2024). It employs two interconnected components: a Subgoal Model that iteratively divides complex tasks into subtasks by predicting the subgoal of the next step (high-level transductive reasoning) and a synthesis model that generates programs to solve these subtasks (low-level inductive reasoning). This structured framework enables the Subgoal Model to acquire broadly applicable skills essential for addressing diverse and complex tasks, while the synthesis model focuses on their implementation, outperforming large Language Models on two program synthesis domains Shi et al. (2023b).
@@ 3.2 REGISM
To further investigate the role of ExeDec’s Subgoal Model, we developed Repeated Execution-Guided Invoking of Synthesis Model (REGISM), a modified version of the ExeDec framework that excludes the Subgoal Model. REGISM invokes the Synthesizer Model to generate a single-step program, i.e., a program consisting of one function only, executes this program, and updates the task specifications using the execution output. At a high level, REGISM shares the same architectural setup and workflow as ExeDec’s No-Subgoal Ablation. The key distinction lies in the training paradigm and the role of the synthesis model.
In both REGISM and the No-Subgoal Ablation, the synthesis model generates a single-step program, executes it, and updates the task specification using the execution results. This iterative process creates new subtask inputs for subsequent steps. The primary difference between the two settings lies in how the models are trained. The No-Subgoal Ablation model is trained to predict subprograms using the ground truth subtask inputs and the task outputs. This forces the model to implicitly infer both decomposition strategies and program generation, making the learning process more entangled. In contrast, the Synthesizer Model used in REGISM is trained to predict subprograms using ground truth subtask inputs and subtask outputs, as it is originally designed for use with ExeDec’s decomposition model. Consequently, the Synthesizer Model is trained to focus purely on program generation without the added complexity of decomposition. During testing, when subtask outputs are unknown, both models receive subtask inputs and task outputs as input. This distinction in training objectives enables us to assess the impact of explicitly supervising decomposition strategies, as done in ExeDec, compared to relying purely on program synthesis, as in REGISM.
@@ 3.3.1 Domains
String Manipulation The RobustFill domain, introduced by Devlin et al. (2017), focuses on synthesizing sequences of string manipulation operations. Given I/O string pairs, the goal is to generate programs that transform inputs into corresponding outputs using a DSL (Appendix B). The DSL includes functions for extracting substrings, modifying strings, and composing operations by applying modifications to prior results. It also supports constant string characters. RobustFill programs are structured as concatenations of largely independent expressions, except for the Compose operation, which applies a function to another’s result. This weak interdependence among subtasks reduces the combinatorial space, simplifying task decomposition learning. However, it also limits error correction during synthesis, making precise decomposition and implementation critical for correct program generation.
@@ 4.1 Domain Differences and Quality of the Subgoal Model
To evaluate ExeDec’s ability to learn task decompositions, we analyze its performance relative to the ground truth. Figure 3 presents a density plot illustrating the distribution of solved tasks.
Figure 3: Tasks solved by ExeDec, clustered based on subtask and subprogram accuracy. The x-axis represents subtask accuracy per task (overlap between predicted and ground truth subtasks), while the y-axis indicates subprogram accuracy (overlap between predicted programs and ground truth solutions). Values are averaged across compositional generalization categories, based on 8,900+ solved tasks in DeepCoder and 26,000+ in RobustFill.
The x-axis represents the subtask accuracy per task, measured as the overlap between predicted subtasks and ground truth subtasks. The y-axis indicates subprogram accuracy, defined as the overlap between predicted programs and the ground truth solutions. Each task’s position in the plot reflects its accuracy: for a task requiring four decomposition steps, if ExeDec correctly predicts two subtask specifications and solves three subtasks with ground-truth-matching programs, it appears at $(x=50\%,y=75\%)$. Consequently, the density plot highlights distinct solution patterns. In the top-right quadrant, both subtasks and subprograms closely align with the ground truth. The bottom-right region represents cases where subtasks mostly match the ground truth, but subprograms differ due to alternative syntactic implementations. The bottom-left quadrant contains instances where both subtasks and subprograms deviate from the ground truth. Finally, in the top-left region, subprograms align closely with the ground truth while subtasks differ. This visualization provides insight into the semantic and syntactic accuracy of ExeDec’s solutions.
@@ 4.2 REGISM as a Driving Force in Boosting Performance
REGISM performs close to ExeDec on average To assess the impact of repeated Synthesizer Model invocation, we evaluate REGISM on the same tasks as ExeDec. As shown in Figure 5, REGISM outperforms the sole Synthesizer Model, highlighting the benefits of execution guidance and repeated invocation. ExeDec achieves higher accuracy on the training distribution and excels in tasks requiring the composition of different concepts, leveraging decomposition to identify and integrate relevant subcomponents. REGISM struggles with length generalization, likely due to its single-step training paradigm, which makes deriving correct subprograms for multi-step tasks more challenging. However, when excluding length generalization tasks, REGISM performs only slightly worse than ExeDec, suggesting that repeated invocation is a key factor in solving tasks. While ExeDec’s decomposition strategies improve performance in some categories, they offer limited or no advantage in others and can even hinder performance in certain cases. The comparable performance of REGISM and ExeDec indicates that task-solving effectiveness primarily stems from repeated invocation, independent of explicit or implicit decomposition guidance.
@@
Comparing REGISM’s results to ExeDec (Figure 4), the tasks solved by REGISM proof to be of similar difficulty to those solved by ExeDec because the number of ground truth decompositions is comparable. Moreover, REGISM solves tasks in about as many iterations as decompositions are intended by the ground truth, i.e., in significantly fewer steps than ExeDec. This indicates that REGISM solves tasks closer to the ground truth solution thus raising the hypothesis that REGISM’s inherent decompositions are closer to the ground truth decomposition than the explicit ones of ExeDec. To investigate this, we plot the subprogram accuracy against the overlap with the ground truth decomposition after executing the subprograms generated by REGISM (Figure 7). This analysis is further split by tasks solved by both approaches or exclusively by either approach. Our findings reveal several notable patterns: REGISM’s task solutions tend to spread along the main diagonal. There is one prominent cluster at the right edge of each plot, where all subtasks align with the ground truth decomposition. To a similar but lesser extent than in ExeDec, REGISM’s largest cluster is located in the bottom-left quadrant. However, REGISM’s solutions exhibit a more uniform distribution along the x-axis compared to ExeDec, whose distribution is distinctly shifted toward lower subtask accuracies. This effect is slightly more pronounced for tasks solved exclusively by REGISM than for those solved by both approaches.
@@
These findings indicate that the Synthesizer Model within REGISM generates solutions closer to the ground truth than ExeDec despite not being explicitly or implicitly trained for this purpose. However, for both approaches, the quality of decompositions does not differ significantly between tasks solved exclusively and those solved by both methods. This finding leaves open the question of why ExeDec solves certain tasks that REGISM does not, especially when the predicted subgoals in those tasks are not more precise than in tasks solved by REGISM, and vice versa.
@@ 5 Related Work
Programming-by-Example Neural networks play a crucial role in guiding program synthesis for PBE (Balog et al., 2016; Yin & Neubig, 2017; Lee et al., 2018). Many approaches divide synthesis into multiple steps, enabling iterative refinement or task planning. Search-based methods include top-down approaches that recursively expand partial solutions (Nye et al., 2019) and bottom-up methods that solve smaller subproblems before combining them into complete programs (Shi et al., 2022; 2023a; Odena et al., 2020). Planning-based techniques first generate abstract program structures, which are later refined into executable code (Murali et al., 2017; Nye et al., 2019; Chen et al., 2020; Hong et al., 2021; Klinger et al., 2023; Prasad et al., 2023; Zhang et al., 2023; Cano et al., 2023). Execution-guided approaches iteratively refine programs using intermediate feedback, such as partial executions or state traces (Ellis et al., 2019; Chen et al., 2018; Shrivastava et al., 2021). Recent advances in Large Language Models have further improved program synthesis by leveraging pre-training on large-scale corpora of code and natural language (Li & Ellis, 2024; Li et al., 2024; Tang et al., 2024). These models enhance synthesis frameworks by capturing semantic and structural patterns of code, aiding in both program generation and refinement.
Decomposition-based Program Synthesis Multi-step synthesis methods inherently break tasks into subtasks as part of their iterative process. In contrast, decomposition approaches explicitly partition a task before leveraging this structure to improve synthesis. Some methods model relationships between subprograms, enabling relational reasoning to identify dependencies and shared structures across subtasks (Hocquette & Cropper, 2023; Hocquette et al., 2024; Hocquette & Cropper, 2024). Similarly, hierarchical composition techniques systematically construct programs by progressively assembling higher-level abstractions from lower-level components (Liu et al., 2023).
@@ 6 Conclusion
Hybrid approaches that integrate transductive and inductive strategies have shown promise in improving program synthesis (Li et al., 2024). ExeDec exemplifies such a hybrid model, where the Subgoal Model represents the transductive component and the Synthesizer Model embodies the inductive aspect, working together to enable compositional generalization. Our results show that both REGISM and ExeDec successfully solve a diverse range of tasks, suggesting potential benefits of framing ExeDec as a hybrid approach. However, identifying the specific conditions under which this combination is most effective remains an open question for future research.
Acknowledgments
Janis Zenkner is supported by the German Federal Ministry for Digital and Transport (BMDV) and the German Federal Ministry for Economic Affairs and Climate Action (BMWK).